40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 17 02:57:32 2014, Andres, Update, 40m Xend Table upgrade, FInish Calculation on Current X-arm mode Matching RawDataForTheModeGreenScan.pngResultForModeMatching.pngDataAndCalculationOfModeMismatch.zip
    Reply  Fri Jul 18 16:52:56 2014, Andres, Update, 40m Xend Table upgrade, FInish Calculation on Current X-arm mode Matching OldAndNewSetupPlotsOfDisplacementAndAngleAtTheETM.pngOldSetUpDisplacementAndNewSetup.m.zip
Message ID: 10237     Entry time: Fri Jul 18 16:52:56 2014     In reply to: 10226
Author: Andres 
Type: Update 
Category: 40m Xend Table upgrade 
Subject: FInish Calculation on Current X-arm mode Matching 

Quote:

Data and Calculation for the Xarm Current Mode Matching

Two days ago, Nick, Jenne, and I took a measurement for the Green Transmission for the X-arm. I took the data and I analyzed it. The first figure attached below is the raw data plotted. I used the function findpeaks in Matlab, and I found all the peaks. Then, by taking close look at the plot, I chose two peaks as shown in the second figure attached below. I took the ratio of the TEM00 and the High order mode, and I average them. This gave me a Mode Matching of 0.9215, which this value is pretty close to the value that I predicted by using a la Mode in http://nodus.ligo.caltech.edu:8080/40m/10191, which is 0.9343. Nick and I measured the reflected power when the cavity is unlocked and when the cavity is locked, so we measured the PreflUnLocked=52+1µW and PreflOnLocked=16+2µW and the backgroundNoise=0.761µW. Using this information we calculated  Prefl/Pin=0.297. Now, since Prefl/Pin=|Eref/Ein|2, we looked at the electric fields component by using the reflectivity of the mirror we calculated 0.67. The number doesn't agree, but this is because we didn't take into account the losses when making this calculation. I'm working in the calculation that will include the losses.

Today, Nick and I ordered the lenses and the mirrors. I'm working in putting together a representation of how much improvement the new design will give us in comparison to the current setup.

We want to be able to graphically see how much better it is the new optical table setup in comparison to the current optical table setup. In other words, we want to be able to see how displacement of the beam and how much angle change can be obtained at the ETM from changing the mirrors angles independently. Depending on the spread of the mirrors' vectors we can observe whether the Gouy phase is good. In the plot below, the dotted lines correspond to the current set up, and we can see that the lines are not spread from each other, which essentially mean that changing the angles of the two mirrors just contribute to small change in angle and in the displacement of the beam at the ETM, and therefore the Gouy phase is not good. Now on the other hand. The other solid lines correspond to the new setup mirrors. We can observe that the spread of the line of mirror 1 and mirror 4 is almost 90 degrees, which just implies that there is a good Gouy phase different between these two mirrors. For the angles chosen in the plot, I looked at how much the PZT yaw the mirrors from the elog http://nodus.ligo.caltech.edu:8080/40m/8912. In this elog, they give a plot in mrad/v for the pitch and yaw, so I took the range that the PZT can yaw the mirrors, and I converted into mdegrees/v and then I plotted as shown below. I plot for the current setup and for the new setup in the same plot. The matlab code is also attached below.

Attachment 1: OldAndNewSetupPlotsOfDisplacementAndAngleAtTheETM.png  41 kB  | Show | Show all
Attachment 2: OldSetUpDisplacementAndNewSetup.m.zip  1 kB
ELOG V3.1.3-